# Representing Anatomical Trees by Denoising Diffusion of Implicit Neural Fields

### v2

@@ Abstract
Anatomical trees play a central role in clinical diagnosis and treatment planning. However, accurately representing anatomical trees is challenging due to their varying and complex topology and geometry. Traditional methods for representing tree structures, captured using medical imaging, while invaluable for visualizing vascular and bronchial networks, exhibit drawbacks in terms of limited resolution, flexibility, and efficiency. Recently, implicit neural representations (INRs) have emerged as a powerful tool for representing shapes accurately and efficiently. We propose a novel approach for representing anatomical trees using INR, while also capturing the distribution of a set of trees via denoising diffusion in the space of INRs. We accurately capture the intricate geometries and topologies of anatomical trees at any desired resolution. Through extensive qualitative and quantitative evaluation, we demonstrate high-fidelity tree reconstruction with arbitrary resolution yet compact storage, and versatility across anatomical sites and tree complexities. The code will be available here [^1] https://github.com/sinashish/TreeDiffusion.
@@ 1 Introduction
Figure 2: Overview: (a) An INR is optimized for each sample in the training dataset, and then flattened to a 1D vector. (b) During inference, the INR is recovered from the flattened vector, followed by MC to extract the underlying signal. (c) The diffusion transformer takes the flattened vectors as input to model the diffusion process. After training, novel INRs can be sampled and used for downstream tasks via (b). (d) Novel tree structures visualized as mesh.
@@
Our contributions
To address the shortcomings of previous works, we are the first to make the following contributions: (1) We train INRs to achieve a faithful representation of complex anatomical trees and demonstrate their usage in segmenting trees from medical images; (2) we employ diffusion models on the space of INR-represented trees for learning tree distributions and generating plausible novel trees with complex topology; (3) we showcase the versatility of our approach in representing trees of varying dimensionality, complexity, and anatomy; and (4) We quantitatively assess our method’s compactness representation and reconstruction accuracy, at arbitrarily high resolution.
@@ 2 Methodology
We propose a diffusion-based generative approach comprising two steps. First, we optimize an INR realized as an MLP by overfitting its parameters to each tree in a dataset so that each INR faithfully represents each tree as a neural occupancy field (Fig. 2(a)). Second, the optimized INRs are flattened into 1D vectors and used to train a denoising diffusion model (DDM) that captures the statistical distribution of the trees (Fig. 2(c)). The DDM is used to synthesize new INRs, via a reverse diffusion process (Fig. 2(b)), which represents the occupancy field (or a mesh, if marching cubes [17] is applied) of the tree (Fig. 2(d)).
@@ 4 Conclusion
We presented the first work to use implicit neural fields for faithful representation of topologically-complex anatomical trees and learnt their distribution via training a diffusion model in the space of neural fields. We demonstrated quantitatively and qualitatively the advantages of our method: versatility (e.g., 2D/3D; vascular/airway; simple/complex topology), lower memory footprint while achieving highly accurate reconstructions at arbitrary high resolutions; synthesis of plausible trees; and application to segmentation of medical images. Further, our representation is amenable to integration into deep learning pipelines and can be easily transformed into other shape representations. Future work may include integrating our method into more advanced deep segmentation pipelines while encoding semantic annotations of tree parts.

### v3

# TrIND: Representing Anatomical Trees by Denoising Diffusion of Implicit Neural Fields

@@ Abstract
Anatomical trees play a central role in clinical diagnosis and treatment planning. However, accurately representing anatomical trees is challenging due to their varying and complex topology and geometry. Traditional methods for representing tree structures, captured using medical imaging, while invaluable for visualizing vascular and bronchial networks, exhibit drawbacks in terms of limited resolution, flexibility, and efficiency. Recently, implicit neural representations (INRs) have emerged as a powerful tool for representing shapes accurately and efficiently. We propose a novel approach, TrIND, for representing anatomical trees using INR, while also capturing the distribution of a set of trees via denoising diffusion in the space of INRs. We accurately capture the intricate geometries and topologies of anatomical trees at any desired resolution. Through extensive qualitative and quantitative evaluation, we demonstrate high-fidelity tree reconstruction with arbitrary resolution yet compact storage, and versatility across anatomical sites and tree complexities. The code is available at: https://github.com/sinashish/TreeDiffusion.
@@ 1 Introduction
Figure 2: TrIND overview: (a) An INR is optimized for each sample in the training dataset, and then flattened to a 1D vector. (b) During inference, the INR is recovered from the flattened vector, followed by MC to extract the underlying signal. (c) The diffusion transformer takes the flattened vectors as input to model the diffusion process. After training, novel INRs can be sampled and used for downstream tasks via (b). (d) Novel tree structures visualized as mesh.
@@
Our contributions.
To address the shortcomings of previous works, we are the first to make the following contributions: (1) We train INRs to achieve a faithful representation of complex anatomical trees and demonstrate their usage in segmenting trees from medical images; (2) we employ diffusion models on the space of INR-represented trees for learning tree distributions and generating plausible novel trees with complex topology; (3) we showcase the versatility of TrIND in representing trees of varying dimensionality, complexity, and anatomy; and (4) We quantitatively assess our method’s compactness representation and reconstruction accuracy, at arbitrarily high resolution.
@@ 2 Methodology
We propose TrIND (pronounced as Trendy), a diffusion-based generative approach comprising two steps. First, we optimize an INR realized as an MLP by overfitting its parameters to each tree in a dataset so that each INR faithfully represents each tree as a neural occupancy field (Fig. 2(a)). Second, the optimized INRs are flattened into 1D vectors and used to train a denoising diffusion model (DDM) that captures the statistical distribution of the trees (Fig. 2(c)). The DDM is used to synthesize new INRs, via a reverse diffusion process (Fig. 2(b)), which represents the occupancy field (or a mesh, if marching cubes [17] is applied) of the tree (Fig. 2(d)).
@@ 4 Conclusion
We presented TrIND, the first work to use implicit neural fields for faithful representation of topologically-complex anatomical trees and learnt their distribution via training a diffusion model in the space of neural fields. We demonstrated quantitatively and qualitatively the advantages of our method: versatility (e.g., 2D/3D; vascular/airway; simple/complex topology), lower memory footprint while achieving highly accurate reconstructions at arbitrary high resolutions; synthesis of plausible trees; and application to segmentation of medical images. Further, our representation is amenable to integration into deep learning pipelines and can be easily transformed into other shape representations. Future work may include integrating our method into more advanced deep segmentation pipelines while encoding semantic annotations of tree parts.